# references.bib

### v2

@@ 3.4 Reinforcement learning
To enhance the model’s reasoning capabilities in audio understanding and speech interaction, we implement a multi-stage reinforcement learning strategy. We leverage our reasoning-centric datasets from SFT and utilize two stages of proximal policy optimization (PPO) [rafailov2023direct] to optimize reasoning efficiency for real-time audio engagement. In the first stage, a binary reward function is employed to limit the thinking sequence length to a predefined maximum. This reward function assigns a value of 1 for reasoning that is appropriately concise (neither empty nor excessively long) and 0 otherwise. Training is conducted for 60 iterations with a global batch size of 64, using an actor learning rate of $1\times 10^{-6}$ and a critic learning rate of $2.5\times 10^{-6}$. The second stage transitions from binary rewards to learned preference scoring, utilizing a trained reward model to evaluate response quality. This stage involves an additional 120 iterations while maintaining the same batch size and learning rate settings. Finally, we incorporate group relative policy optimization (GRPO) [rafailov2023direct] for 400 iterations to further improve the model’s audio perceptual abilities.
@@ 4.1 Automatic speech recognition
As the most critical component of audio understanding and speech interaction, we first evaluate the model’s capability in automatic speech recognition. We evaluate Step-Audio 2 across six Chinese test sets, four English test sets, three multilingual test sets (Japanese, Cantonese, Arabic), and six in-house Chinese dialect and accented Mandarin test sets. For comparative analysis, we utilize top-performing models from both open-source and commercial domains as baselines, including Doubao LLM ASR [^1] Doubao LLM ASR refers to https://www.volcengine.com/docs/6561/1354868, GPT-4o Transcribe [^2] GPT-4o Transcribe is evaluated using its latest model, gpt-4o-transcribe, via its official API., Kimi-Audio [ding2025kimi], and Qwen-Omni. We prefer GPT-4o Transcribe than GPT-4o Audio since the formers provide stronger results. Notably, Doubao LLM ASR and GPT-4o Transcribe represent specialized ASR systems that achieve leading-edge performance.
We evaluate all the models without specifying language [^3] We evaluate without specifying language to ensure a fair comparison. Notably, Qwen-Omni lacks a language-independent testing approach, specifying language may yield better results. and summarize the results in Table 1. Step-Audio 2 outperforms existing open-source and commercial ASR models in both general English and Chinese recognition, achieving an average word error rate (WER) of 3.18% on English and an average character error rate (CER) of 3.11% on Chinese test sets. Moreover, Step-Audio 2 offers comparable results to GPT-4o Transcribe on Arabian and Japanese recognition, to Qwen-Omni on Cantonese recognition, demonstrating its capability in multilingual speech recognition. In addition, Step-Audio 2 achieves the lowest average CER among 4 in-house Chinese accented Mandarin and 2 dialect test sets. These results highlight the superiority of Step-Audio 2 in understanding the semantic information in speech.
@@
English Common Voice 9.20 9.30 7.83 8.33 5.98
FLEURS English 7.22 2.71 4.47 5.05 3.05
LibriSpeech clean 2.92 1.75 1.49 2.93 1.19
LibriSpeech other 5.32 4.23 2.91 5.07 2.49
Average 6.17 4.50 4.18 5.35 3.18
Chinese AISHELL 0.98 3.52 0.64 1.17 0.65
AISHELL-2 3.10 4.26 2.67 2.40 2.13
FLEURS Chinese 2.92 2.62 2.91 7.01 2.80
KeSpeech phase1 6.48 26.80 5.11 6.45 3.62
WenetSpeech meeting 4.90 31.40 5.21 6.61 4.73
WenetSpeech net 4.46 15.71 5.93 5.24 4.74
Average 3.81 14.05 3.75 4.81 3.11
FLEURS Arabian N/A 11.72 N/A 25.13 15.66
Multilingual Common Voice yue 9.20 11.10 38.90 7.89 8.04
FLEURS Japanese N/A 3.27 N/A 10.49 3.44
Anhui accent 8.83 50.55 22.17 18.73 10.99
Guangdong accent 4.99 7.83 3.76 4.03 3.87
Guangxi accent 3.37 7.09 4.29 3.35 4.08
In-house Shanxi accent 20.26 55.03 34.71 25.95 13.77
Sichuan dialect 3.01 32.85 5.26 5.61 4.28
Shanghai dialect 47.49 89.58 82.90 58.74 18.14
Average 14.66 40.49 25.52 19.40 9.19
@@ 4.2 Paralinguistic information understanding
We evaluate GPT-4o Audio, Kimi-Audio, Qwen-Omni, Step-Audio-AQAA, and Step-Audio 2 using the StepEval-Audio-Paralinguistic benchmark, with results presented in Table 2. The experimental results highlight the comprehensive capabilities of Step-Audio 2 in understanding various paralinguistic information, achieving an average accuracy of 76.55, which is a significant improvement over other baseline models.
@@ 4.3 Audio understanding
We then assess Step-Audio 2’s general audio comprehension across sound, speech, and music using the latest version of the MMAU benchmark [sakshi2024mmau] [^4] MMAU v05.15.25 test-mini.
As baselines, we employ Audio Flamingo 3, Gemini 2.5 Pro, GPT-4o Audio, Kimi-Audio, Omni-R1 [rouditchenko2025omni], Qwen2.5-Omni, and Step-Audio-AQAA. We obtain the reported results for Audio Flamingo 3, Omni-R1, and Qwen2.5-Omni from their original papers. The results of Gemini 2.5 Pro are obtained from the official website of MMAU. And we re-evaluate GPT-4o Audio, Kimi-Audio and Step-Audio-AQAA due to the recent update of the MMAU benchmark.
The results are summarized in Table 3. Step-Audio 2 achieves the highest average score of 77.4%, followed by Omni-R1 and Audio Flamingo 3, both of which are specialized approaches in audio understanding. Specifically, Step-Audio 2 yields the best results in sound and music tracks and on par results with the best in speech track, demonstrating its versatility and robustness across different audio domains.
@@ 4.4 Speech translation
We evaluate the model’s bidirectional Chinese-English speech translation capabilities using two benchmarks: speech-to-text translation (S2TT) on CoVoST 2 [wang2020covost2massivelymultilingual] and speech-to-speech translation (S2ST) on CVSS [jia2022cvsscorpusmassivelymultilingual]. Additionally, we use the reported results of Qwen2.5-Omni for CoVoST 2, while for CVSS, we employ Qwen-Omni as a baseline. Kimi-Audio is excluded because it consistently ignores prompts and performs ASR instead of translation. Using BLEU as the evaluation metric, the results in Table 4 demonstrate that Step-Audio 2 achieves superior performance in Chinese-English bidirectional translations, obtaining the highest average score on both the CoVoST 2 and CVSS test sets.
@@ 4.6 Speech-to-speech conversation
We finally employ URO-Bench [yan2025urobenchcomprehensivebenchmarkendtoend] to evaluate Step-Audio 2 and other open-source and commercial LALMs, including GPT-4o Audio, Kimi-Audio, Qwen-Omni, and Step-Audio-AQAA. URO-Bench consists of 16 and 20 datasets on two difficulty tracks, evaluating the model’s understanding, reasoning and oral conversation abilities, such as ASR, instruction following, commonsense knowledge, mathematics, and speech naturalness, emotion and speaking styles expressions. We follow the ASR-mediated procedure in URO-Bench for evaluation, employing Whisper for ASR and GPT-4o-mini for automatic judging.
As demonstrated in Table 6, Step-Audio 2 significantly outperforms existing large audio language models, including GPT-4o Audio, in Chinese speech-to-speech conversation scenarios, achieving the highest average scores of 78.86 on the basic track and 70.83 on the pro track. In English speech-to-speech conversations, while Step-Audio 2 is slightly outperformed by GPT-4o Audio, it provides very competitive results and exceeds the other approaches. More detailed results on URO-Bench are provided in Table 7.
@@
GPT-4o Audio Chinese 74.18 82.98 57.23 82.33 66.91 72.94 51.52 71.14
Kimi-Audio 70.47 75.86 59.69 75.85 66.21 63.13 55.09 76.70
Qwen-Omni 62.08 46.44 64.73 75.05 61.06 61.55 59.79 61.43
Step-Audio-AQAA 55.73 66.02 57.31 43.87 59.15 61.82 52.74 60.74
Step-Audio 2 78.86 87.66 68.52 80.39 70.83 79.35 59.71 69.72
@@ A.4 External contributors
Appendix B Detailed evaluation results on URO-bench
Table 7: Detailed results on the URO-Bench. C. E. stand for Chinese and English. B. P. stand for the basic and pro tracks. U. R. O. stand for understanding, reasoning and oral conversation, respectively.
Model C. B. U. C. B. U. C. B. R. C. B. R. C. B. O. C. B. O.
Repeat LCSTS MLC OpenBook Alpaca Claude
GPT-4o Audio 86.11 79.85 50.00 64.46 75.87 88.79
Kimi-Audio 73.32 78.40 52.45 66.93 69.90 81.81
Qwen-Omni 19.66 73.22 61.76 67.70 70.40 79.71
Step-Audio-AQAA 63.72 68.33 67.40 47.21 56.23 31.50
Step-Audio 2 96.16 79.16 62.99 74.06 76.23 84.54
Model C. P. U. C. P. U. C. P. U. C. P. R. C. P. R. C. P. O.
UnderEmotion CodeSwitch Safety MLC-Pro Speaker SRT
GPT-4o Audio 72.07 63.43 83.33 53.65 49.39 85.71
Kimi-Audio 76.29 66.76 46.33 58.33 51.84 99.05
Qwen-Omni 70.13 66.86 47.67 72.92 46.67 98.10
Step-Audio-AQAA 70.46 52.00 63.00 53.65 51.84 97.14
Step-Audio 2 78.14 63.90 96.00 70.31 49.12 96.19
Model C. P. O. C. P. O. E. B. U. E. B. U. E. B. U. E. B. R.
GenEmotion GenStyle Repeat Summary Gaokao Storal
GPT-4o Audio 30.28 97.44 96.08 92.94 81.52 83.15
Kimi-Audio 35.49 95.56 90.27 78.19 81.63 70.88
Qwen-Omni 8.58 77.61 28.36 80.51 89.99 72.90
Step-Audio-AQAA 26.80 58.29 93.02 90.62 86.80 53.00
Step-Audio 2 23.56 89.40 97.74 83.90 90.76 76.38
Model E. B. R. E. B. R. E. B. R. E. B. O. E. B. O. E. B. O.
Truthful Gsm8k MLC Alpaca Common Wildchat
GPT-4o Audio 74.27 70.10 76.08 94.00 85.73 91.50
Kimi-Audio 53.90 0.97 43.50 63.05 51.40 66.63
Qwen-Omni 62.03 63.52 80.04 78.16 72.40 77.92
Step-Audio-AQAA 62.30 38.20 71.00 71.86 75.53 68.79
Step-Audio 2 65.45 62.83 77.02 81.74 72.87 81.62
Model E. P. U. E. P. U. E. P. U. E. P. U. E. P. U. E. P. R.
UnderEmotion CodeSwitch Safety Clotho MuCho MLC-Pro
GPT-4o Audio 54.21 72.00 88.61 42.77 45.66 71.43
Kimi-Audio 56.40 62.86 57.78 43.27 31.30 23.81
Qwen-Omni 52.21 61.52 63.06 13.58 32.15 66.30
Step-Audio-AQAA 48.86 68.95 63.06 8.43 31.94 45.05
Step-Audio 2 58.05 56.48 91.11 37.74 58.95 69.60
Model E. P. R. E. P. R. E. P. O. E. P. O. E. P. O. E. P. O.
MtBench Speaker SRT GenEmotion GenStyle Multilingual
GPT-4o Audio 76.07 45.58 80.93 37.62 100.00 95.28
Kimi-Audio 55.54 42.42 89.30 25.33 80.61 28.90
Qwen-Omni 74.81 50.55 86.51 8.89 80.30 21.94
Step-Audio-AQAA 68.74 49.82 90.70 29.62 82.12 36.81
Step-Audio 2 68.46 45.58 90.70 27.70 81.82 36.77

### v3

@@ 3.4 Reinforcement learning
To enhance the model’s reasoning capabilities in audio understanding and speech interaction, we implement a multi-stage reinforcement learning strategy. We leverage our reasoning-centric datasets from SFT and utilize two stages of proximal policy optimization (PPO) [54] to optimize reasoning efficiency for real-time audio engagement. In the first stage, a binary reward function is employed to limit the thinking sequence length to a predefined maximum. This reward function assigns a value of 1 for reasoning that is appropriately concise (neither empty nor excessively long) and 0 otherwise. Training is conducted for 60 iterations with a global batch size of 64, using an actor learning rate of $1\times 10^{-6}$ and a critic learning rate of $2.5\times 10^{-6}$. The second stage transitions from binary rewards to learned preference scoring, utilizing a trained reward model to evaluate response quality. This stage involves an additional 120 iterations while maintaining the same batch size and learning rate settings. Finally, we incorporate group relative policy optimization (GRPO) [54] for 400 iterations to further improve the model’s audio perceptual abilities.
@@ 4.1 Automatic speech recognition
As the most critical component of audio understanding and speech interaction, we first evaluate the model’s capability in automatic speech recognition. We evaluate Step-Audio 2 across six Chinese test sets, four English test sets, three multilingual test sets (Japanese, Cantonese, Arabic), and six in-house Chinese dialect and accented Mandarin test sets. For comparative analysis, we utilize top-performing models from both open-source and commercial domains as baselines, including Doubao LLM ASR [^1] Doubao LLM ASR refers to https://www.volcengine.com/docs/6561/1354868, GPT-4o Transcribe [^2] GPT-4o Transcribe is evaluated using its latest model, gpt-4o-transcribe, via its official API., Kimi-Audio [18], and Qwen-Omni. We prefer GPT-4o Transcribe than GPT-4o Audio since the formers provide stronger results. Notably, Doubao LLM ASR and GPT-4o Transcribe represent specialized ASR systems that achieve leading-edge performance.
We evaluate all the models without specifying language [^3] We evaluate without specifying language to ensure a fair comparison. Notably, Qwen-Omni lacks a language-independent testing approach, specifying language may yield better results. and summarize the results in Table 1. Step-Audio 2 outperforms existing open-source and commercial ASR models in both general English and Chinese recognition, achieving an average word error rate (WER) of 3.14% on English and an average character error rate (CER) of 3.08% on Chinese test sets. Moreover, Step-Audio 2 offers comparable results to GPT-4o Transcribe on Arabian and Japanese recognition, to Qwen-Omni on Cantonese recognition, demonstrating its capability in multilingual speech recognition. In addition, Step-Audio 2 achieves the lowest average CER among 4 in-house Chinese accented Mandarin and 2 dialect test sets. These results highlight the superiority of Step-Audio 2 in understanding the semantic information in speech.
@@
English Common Voice 9.20 9.30 7.83 8.33 5.95
FLEURS English 7.22 2.71 4.47 5.05 3.03
LibriSpeech clean 2.92 1.75 1.49 2.93 1.17
LibriSpeech other 5.32 4.23 2.91 5.07 2.42
Average 6.17 4.50 4.18 5.35 3.14
Chinese AISHELL 0.98 3.52 0.64 1.17 0.63
AISHELL-2 3.10 4.26 2.67 2.40 2.10
FLEURS Chinese 2.92 2.62 2.91 7.01 2.68
KeSpeech phase1 6.48 26.80 5.11 6.45 3.63
WenetSpeech meeting 4.90 31.40 5.21 6.61 4.75
WenetSpeech net 4.46 15.71 5.93 5.24 4.67
Average 3.81 14.05 3.75 4.81 3.08
FLEURS Arabian N/A 11.72 N/A 25.13 14.22
Multilingual Common Voice yue 9.20 11.10 38.90 7.89 7.90
FLEURS Japanese N/A 3.27 N/A 10.49 3.18
Anhui accent 8.83 50.55 22.17 18.73 10.61
Guangdong accent 4.99 7.83 3.76 4.03 3.81
Guangxi accent 3.37 7.09 4.29 3.35 4.11
In-house Shanxi accent 20.26 55.03 34.71 25.95 12.44
Sichuan dialect 3.01 32.85 5.26 5.61 4.35
Shanghai dialect 47.49 89.58 82.90 58.74 17.77
Average 14.66 40.49 25.52 19.40 8.85
@@ 4.2 Paralinguistic information understanding
We evaluate GPT-4o Audio, Kimi-Audio, Qwen-Omni, Step-Audio-AQAA, and Step-Audio 2 using the StepEval-Audio-Paralinguistic benchmark, with results presented in Table 2. The experimental results highlight the comprehensive capabilities of Step-Audio 2 in understanding various paralinguistic information, achieving an average accuracy of 83.09, which is a significant improvement over other baseline models.
@@ 4.3 Audio understanding
We then assess Step-Audio 2’s general audio comprehension across sound, speech, and music using the latest version of the MMAU benchmark [58] [^4] MMAU v05.15.25 test-mini.
As baselines, we employ Audio Flamingo 3, Gemini 2.5 Pro, GPT-4o Audio, Kimi-Audio, Omni-R1 [56], Qwen2.5-Omni, and Step-Audio-AQAA. We obtain the reported results for Audio Flamingo 3, Omni-R1, and Qwen2.5-Omni from their original papers. The results of Gemini 2.5 Pro are obtained from the official website of MMAU. And we re-evaluate GPT-4o Audio, Kimi-Audio and Step-Audio-AQAA due to the recent update of the MMAU benchmark.
The results are summarized in Table 3. Step-Audio 2 achieves the highest average score of 78.0, followed by Omni-R1 and Audio Flamingo 3, both of which are specialized approaches in audio understanding. Specifically, Step-Audio 2 yields the best results in sound and speech tracks and on par results with the best in music track, demonstrating its versatility and robustness across different audio domains.
@@ 4.4 Speech translation
We evaluate the model’s bidirectional Chinese-English speech translation capabilities using two benchmarks: speech-to-text translation (S2TT) on CoVoST 2 [63] and speech-to-speech translation (S2ST) on CVSS [37]. Additionally, we use the reported results of Qwen2.5-Omni for CoVoST 2, while for CVSS, we employ Qwen-Omni as a baseline. Kimi-Audio is excluded because it consistently ignores prompts and performs ASR instead of translation. Using BLEU as the evaluation metric, the results in Table 4 demonstrate that Step-Audio 2 achieves superior performance in Chinese-English bidirectional translations, obtaining the highest average score on both the CoVoST 2 and CVSS test sets.
@@ 4.6 Speech-to-speech conversation
We finally employ URO-Bench [75] to evaluate Step-Audio 2 and other open-source and commercial LALMs, including GPT-4o Audio, Kimi-Audio, Qwen-Omni, and Step-Audio-AQAA. URO-Bench consists of multiple datasets on two difficulty tracks, evaluating the model’s understanding, reasoning and oral conversation abilities, such as ASR, instruction following, commonsense knowledge, mathematics, and speech naturalness, emotion and speaking styles expressions. We follow the ASR-mediated procedure in URO-Bench for evaluation, employing Whisper for ASR and GPT-4o-mini for automatic judging.
As demonstrated in Table 6, Step-Audio 2 significantly outperforms existing large audio language models, including GPT-4o Audio, in Chinese speech-to-speech conversation scenarios, achieving the highest average scores of 83.32 on the basic track and 68.25 on the pro track. In English speech-to-speech conversations, while Step-Audio 2 is slightly outperformed by GPT-4o Audio, it provides very competitive results and exceeds the other approaches.
@@
GPT-4o Audio Chinese 78.59 89.40 65.48 85.24 67.10 70.60 57.22 70.20
Kimi-Audio 73.59 79.34 64.66 79.75 66.07 60.44 59.29 76.21
Qwen-Omni 68.98 59.66 69.74 77.27 59.11 59.01 59.82 58.74
Step-Audio-AQAA 74.71 87.61 59.63 81.93 65.61 74.76 47.29 68.97
Step-Audio 2 83.32 91.05 75.45 86.08 68.25 74.78 63.18 65.10
@@ A.4 External contributors
Appendix B Introduction and evaluation results of Step-Audio 2 mini
We are pleased to release Step-Audio 2 mini, a special open-source version of Step-Audio 2, available at https://github.com/stepfun-ai/Step-Audio2. Step-Audio 2 mini employs the encoder from Qwen2-Audio as its audio encoder and is initialized with Qwen2.5-7B. Step-Audio 2 mini is trained on the same dataset as Step-Audio 2, but it is limited to utilize only the web search tool.
Step-Audio 2 mini is a more developer-friendly variant of Step-Audio 2, with a parameter count for fair comparisons with open-source models including Qwen-Omni and Kimi-Audio. Evaluation results [^5] Evaluation results are obtained with our vLLM backend and may differ from the results with transformers backend. demonstrate that Step-Audio 2 mini delivers on par results with Step-Audio 2, exceeding most open-source and commercial models such as GPT-4o Audio.
B.1 Automatic speech recognition
Table 7: Comparison between Doubao LLM ASR, GPT-4o Transcribe, Kimi-Audio, Qwen-Omni, Step-Audio 2 and Step-Audio 2 mini, on character (for Chinese, Cantonese and Japanese) and word (for Arabian and English) error rates among multiple ASR test sets. N/A indicates that the language is not supported.
Category Test set Doubao LLM ASR GPT-4o Transcribe Kimi- Audio Qwen- Omni Step- Audio 2 Step-Audio 2 mini
English Common Voice 9.20 9.30 7.83 8.33 5.95 6.76
FLEURS English 7.22 2.71 4.47 5.05 3.03 3.05
LibriSpeech clean 2.92 1.75 1.49 2.93 1.17 1.33
LibriSpeech other 5.32 4.23 2.91 5.07 2.42 2.86
Average 6.17 4.50 4.18 5.35 3.14 3.50
Chinese AISHELL 0.98 3.52 0.64 1.17 0.63 0.78
AISHELL-2 3.10 4.26 2.67 2.40 2.10 2.16
FLEURS Chinese 2.92 2.62 2.91 7.01 2.68 2.53
KeSpeech phase1 6.48 26.80 5.11 6.45 3.63 3.97
WenetSpeech meeting 4.90 31.40 5.21 6.61 4.75 4.87
WenetSpeech net 4.46 15.71 5.93 5.24 4.67 4.82
Average 3.81 14.05 3.75 4.81 3.08 3.19
FLEURS Arabian N/A 11.72 N/A 25.13 14.22 16.46
Multilingual Common Voice yue 9.20 11.10 38.90 7.89 7.90 8.32
FLEURS Japanese N/A 3.27 N/A 10.49 3.18 4.67
Anhui accent 8.83 50.55 22.17 18.73 10.61 11.65
Guangdong accent 4.99 7.83 3.76 4.03 3.81 4.44
Guangxi accent 3.37 7.09 4.29 3.35 4.11 3.51
In-house Shanxi accent 20.26 55.03 34.71 25.95 12.44 15.60
Sichuan dialect 3.01 32.85 5.26 5.61 4.35 4.57
Shanghai dialect 47.49 89.58 82.90 58.74 17.77 19.30
Average 14.66 40.49 25.52 19.40 8.85 9.85
B.2 Paralinguistic information understanding
Table 8: Comparison between GPT-4o Audio, Kimi-Audio, Qwen-Omni, Step-Audio-AQAA, Step-Audio 2 and Step-Audio 2 mini on StepEval-Audio-Paralinguistic.
Model Avg. Gender Age Timbre Scenario Event
GPT-4o Audio 43.45 18 42 34 22 14
Kimi-Audio 49.64 94 50 10 30 48
Qwen-Omni 44.18 40 50 16 28 42
Step-Audio-AQAA 36.91 70 66 18 14 14
Step-Audio 2 83.09 100 96 82 78 60
Step-Audio 2 mini 80.00 100 94 80 78 60
Model Emotion Pitch Rhythm Speed Style Vocal
GPT-4o Audio 82 40 60 58 64 44
Kimi-Audio 66 56 40 44 54 54
Qwen-Omni 76 32 54 50 50 48
Step-Audio-AQAA 40 38 48 54 44 0
Step-Audio 2 86 82 86 88 88 68
Step-Audio 2 mini 82 82 68 74 86 76
B.3 Audio understanding
Table 9: Comparison between Audio Flamingo 3, Gemini 2.5 Pro, GPT-4o Audio, Kimi-Audio, Omni-R1, Qwen2.5-Omni, Step-Audio-AQAA, Step-Audio 2 and Step-Audio 2 mini on MMAU.
Model Avg. Sound Speech Music
Audio Flamingo 3 73.1 76.9 66.1 73.9
Gemini 2.5 Pro 71.6 75.1 71.5 68.3
GPT-4o Audio 58.1 58.0 64.6 51.8
Kimi-Audio 69.6 79.0 65.5 64.4
Omni-R1 77.0 81.7 76.0 73.4
Qwen2.5-Omni 71.5 78.1 70.6 65.9
Step-Audio-AQAA 49.7 50.5 51.4 47.3
Step-Audio 2 78.0 83.5 76.9 73.7
Step-Audio 2 mini 73.2 76.6 71.5 71.6
B.4 Speech translation
Table 10: Comparison of BLEU scores between GPT-4o Audio, Qwen2.5-Omni, Qwen-Omni, Step-Audio-AQAA, Step-Audio 2 and Step-Audio 2 mini on speech-to-text and speech-to-speech translation.
Model CoVoST 2 (Speech-to-text translation)
Avg. English-to-Chinese Chinese-to-English
GPT-4o Audio 29.61 40.20 19.01
Qwen2.5-Omni 35.40 41.40 29.40
Step-Audio-AQAA 28.57 37.71 19.43
Step-Audio 2 39.26 49.01 29.51
Step-Audio 2 mini 39.29 49.12 29.47
Model CVSS (Speech-to-speech translation)
Avg. English-to-Chinese Chinese-to-English
GPT-4o Audio 23.68 20.07 27.29
Qwen-Omni 15.35 8.04 22.66
Step-Audio-AQAA 27.36 30.74 23.98
Step-Audio 2 30.87 34.83 26.92
Step-Audio 2 mini 29.08 32.81 25.35
B.5 Speech-to-speech conversation
Table 11: Comparison between GPT-4o Audio, Kimi-Audio, Qwen-Omni, Step-Audio-AQAA, Step-Audio 2 and Step-Audio 2 mini on the URO-Bench. U. R. O. stands for understanding, reasoning, and oral conversation, respectively.
Model Language Basic Pro
Avg. U. R. O. Avg. U. R. O.
GPT-4o Audio Chinese 78.59 89.40 65.48 85.24 67.10 70.60 57.22 70.20
Kimi-Audio 73.59 79.34 64.66 79.75 66.07 60.44 59.29 76.21
Qwen-Omni 68.98 59.66 69.74 77.27 59.11 59.01 59.82 58.74
Step-Audio-AQAA 74.71 87.61 59.63 81.93 65.61 74.76 47.29 68.97
Step-Audio 2 83.32 91.05 75.45 86.08 68.25 74.78 63.18 65.10
Step-Audio 2 mini 77.81 89.19 64.53 84.12 69.57 76.84 58.90 69.42
GPT-4o Audio English 84.54 90.18 75.90 90.41 67.51 60.65 64.36 78.46
Kimi-Audio 60.04 83.36 42.31 60.36 49.79 50.32 40.59 56.04
Qwen-Omni 70.58 66.29 69.62 76.16 50.99 44.51 63.88 49.41
Step-Audio-AQAA 71.11 90.15 56.12 72.06 52.01 44.25 54.54 59.81
Step-Audio 2 83.90 92.72 76.51 84.92 66.07 64.86 67.75 66.33
Step-Audio 2 mini 74.36 90.07 60.12 77.65 61.25 58.79 61.94 63.80